# PartGS: Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

### v2

@@ Abstract
Low-level 3D representations, such as point clouds, meshes, NeRFs, and 3D Gaussians, are commonly used to represent 3D objects or scenes. However, human perception typically understands 3D objects at a higher level as a composition of parts or structures rather than points or voxels. Representing 3D objects or scenes as semantic parts can benefit further understanding and applications. In this paper, we introduce PartGS, part-aware 3D reconstruction by a hybrid representation of 2D Gaussians and Superquadrics, which parses objects or scenes into semantic parts, digging 3D structural clues from multi-view image inputs. Accurate structured geometry reconstruction and high-quality rendering are achieved at the same time. Our method simultaneously optimizes superquadric meshes and Gaussians by coupling their parameters within our hybrid representation. On one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes. On the other hand, 2D Gaussians capture complex texture and geometry details, ensuring high-quality appearance and geometry reconstruction. Our method is fully unsupervised and outperforms existing state-of-the-art approaches in extensive experiments on DTU, ShapeNet, and real-life datasets.
@@ 1 Introduction
Figure 1: We propose a pipeline to reconstruct the scene in a part-aware manner. Prior works [22, 27] model the scene using primitives, while the proposed method can further model precise geometry details and textures. Here, EMS [22] takes point cloud inputs and reconstructs non-textured primitives.
@@
3D reconstruction from multi-view images is a long-studying problem in 3D vision and computer graphics. Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which differ from human perception. Humans understand 3D scenes as different semantic parts [26]. For instance, when observing a scene, we mentally construct high-level structural information, such as scene graphs, rather than point clouds or detailed geometry. Thus, we propose to learn part-aware scene reconstruction to decompose scenes into different individual semantic shapes or parts, facilitating tasks including physical simulation, manipulation/editing, 3D content generation, etc.
There are prior methods [38, 51, 2, 28, 20, 44, 22, 23, 36, 42, 48] working on this topic, while they mainly learn the goal with 3D supervisions and cannot preserve accurate geometries, which leaves inconvenience to apply in realistic scenarios. While NeRF [25] has demonstrated great potential in reconstructing textured 3D scenes from multi-view images, some approaches [37, 49] try to learn part-aware objects via NeRFs. In particular, Part-NeRF [37] represents an object as multiple neural radiance fields. However, the complex composition of multiple implicit fields increases learning difficulty, resulting in subpar rendering quality and insufficient semantic decomposition. Another state-of-the-art approach DBW [27] decomposes scenes into blocks represented by superquadric primitives [2], optimizing their parameters and UV texture maps by minimizing rendering loss. Although DBW effectively decomposes 3D scenes into reasonable parts, learning accurate part-aware geometry and appearance still proves challenging as shown in Fig. 1.
@@
To tackle the problem, we propose PartGS, a part-aware hybrid representation that integrates 2D Gaussians [16] and superquadrics [2] to reconstruct scenes that are both well-textured and geometrically accurate. There are previous approaches [15, 13, 39] combining mesh reconstruction and Gaussian splatting, where they first reconstruct the mesh and then attach Gaussians to it. Their purpose is to get both representations, while our goal is to get part segmentation and reconstruction at the same time. Our method directly learns superquadric mesh and Gaussians simultaneously by coupling parameters between them while constraining each part to be represented as a single superquadric. The rendering capability of coupled Gaussians drives the optimization of hybrid representation, enabling the use of their convexity to assist in the decomposition of shapes. The unsupervised training of part-aware reconstruction is built upon the assumption that each functional part of most objects should be a basic shape and can be represented by a superquadric. On the other hand, Gaussian splatting, renowned for its superior rendering capabilities and efficient training[18, 11, 24], is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality significantly. In the hybrid representation, 2D Gaussians are attached to the surface of superquadrics, with their pose and shape parameters coupled to the corresponding superquadric mesh. Compared to the vanilla Gaussian [18] that fills up the whole occupancy with individual-parameter 3D Gaussians, our method is more compact and efficient. The optimized parameters include global ones controlling the overall shape, pose, and opacity of hybrid blocks, and local parameters for spherical harmonic coefficients of the Gaussians.
@@
During training, the parameters are optimized by rendering loss without additional supervision. We observe that image rendering loss tends to settle into local minima during the optimization of superquadrics. To address this, we design several regularizers to enforce global consistency between the 3D representation and the input 2D information. This allows us to segment parts in a fully unsupervised manner, achieving block-level reconstruction. At last, to model the accurate geometry, we conduct a point-level refinement, freeing 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes. Contributions are summarized as follows:
  • A novel hybrid representation is introduced to learn part-aware textured 3D scenes. The hybrid representation combines the benefits of both superquadrics and Gaussian splitting, enabling accurate part reconstruction and efficient high-quality rendering.
  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with superquadric meshes and 2D Gaussians optimized simultaneously by novel regularizers.
  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, enabling applications of part augmentation and physical interactions.
@@
Figure 2: Overview of PartGS. PartGS takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians. The representation initializes from random superquadrics and is gradually optimized during training to get a block-level reconstruction. Then, we free the constraints of Gaussians to model detailed geometry to get a point-level reconstruction. The last column demonstrates one application of PartGS: physics simulations, where one part is picked up from the scene and thrown forward. Please refer to the supplementary materials for more video demos.
2 Related Works
2.1 Primitive-based 3D Representation
Learning structured shape representations aims to decompose objects or scenes into semantically coherent geometric primitives, providing a crucial pathway to shape understanding and generation [43, 50, 14, 9, 41, 35]. The earliest classical approaches, dating back to the last century, such as Blocks World [32] and Generalized Cylinders [3] emphasize compactness as the central goal. A substantial amount of prior work has focused on processing 3D data inputs, such as point clouds, meshes, and voxel grids. These methods decompose the 3D data into various types of primitive ensembles, including cuboids [38, 31], superquadrics [29, 30, 44, 22], and convex shapes [8, 10]. MonteBoxFinder [31], uses a RANSAC-based [12] with cuboids to represent point clouds, combined with Monte Carlo Tree Search for optimal selection. Alternatively, EMS [22] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover primitives. These methods rely on perfect 3D data, which is rarely available due to incompleteness or noise.
@@ 2.1 Primitive-based 3D Representation
Few studies [37, 27, 49, 1] have attempted to create structure-aware 3D representations directly from images. PartNeRF [37] uses ellipses with neural radiance fields, but this spatial-based decomposition limits its ability to capture meaningful structural information. ISCO [1] and DBW [27] use 3D superquadrics for shape decomposition, which enables more meaningful structural separation. However, their simple shape parameters struggle to capture complex geometries, leading to poor geometry and appearance reconstruction. DPA-Net [49] has advanced 3D shape abstraction from sparse views but generates redundant parts and struggles with realistic texture rendering. Our approach introduces a hybrid representation that learns part-aware 3D scenes while ensuring accurate structural reconstruction and efficient, high-quality rendering.
@@ 2.2 Mesh-based Gaussian Splatting
Gaussian splatting(GS) [18] has been rapidly adopted in multiple fields due to its remarkable rendering capability. Several studies [15, 7, 13, 39] attempt to align Gaussians with mesh surfaces for easier editing and animation. Among them, SuGaR [15] uses flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and actual Gaussians. GaMeS [39] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to triangular facets of the mesh. Similarly, Gao et al. [13] proposes a mesh-based GS to achieve large-scale deformation effects on objects. Recently, 2DGS [16] proposes 2D Gaussians for surface modeling and significantly enhances the geometric quality. Inspired by them, we propose a representation that combines 2D Gaussians with superquadrics. A key distinction between our method and previous mesh-based GS approaches is that these methods require first reconstructing the scene’s mesh and then binding Gaussians to the mesh surface, which results in a non-continuous mesh. In contrast, our method directly optimizes the mesh through a rendering loss, enabling part-aware mesh reconstruction.
@@ 3 Method
Given a set of calibrated multi-view images, our goal is to learn part-aware 3D geometry and appearance. A part-aware reconstruction is often represented as a composition of multiple basic shapes (geometric primitives). We propose a primitive-based hybrid representation in Sec. 3.1. In Sec. 3.2, we demonstrate how this hybrid representation is optimized to fit primitives as parts (block-level). In order to further detail the geometry of each part, the parameters of Gaussians are further refined to get a point-level reconstruction.
3.1 Hybrid Representation
As shown in Fig. 2, the proposed hybrid representation combines superquadrics with 2D Gaussians in a compact form. To leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes. This representation retains the superquadric’s ability to effectively parse a 3D scene into distinct parts. Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of previous work [22, 31, 27, 48]. Sharing pose parameters between superquadrics and 2D Gaussians further enhances the representation’s efficiency.
The proposed hybrid model is parametric, controlled by both primitive and Gaussian parameters, which are optimized simultaneously. Considering a 3D scene $\mathcal{S}$, we parse it into multiple superquadrics, i.e. blocks. Each block is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$. Here, $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks. The hybrid blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture. These parameters are optimized via differentiable rendering to parse the target scene.
Shape and Scale Parameters. For each hybrid block $\mathcal{B}_{i}$, its geometry is controlled by superquadric parameters [2]. More concretely, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define the superquadric mesh, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes. Similar to the positioning of vertexes in an icosphere, the 3D position of a superquadric vertex can be obtained by:
@@ 3.1 Hybrid Representation
where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate, respectively. The shape and scale parameters are optimized during training, thus controlling the block deformation to learn part-aware geometry.
Pose Parameters. The pose of the $i$-th hybrid block is defined by its rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. The vertex position from Eq. 1 is transformed from the local coordinate to world space as: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $j$ indexes the vertices. In prior methods such as SuGaR [15] and GaMeS [39], Gaussians are placed on the reconstructed mesh, and they have individual poses and shapes for better appearance modeling. Differently, in our method, the role of Gaussians is to construct a differentiable rendering that bridges images and superquadrics. To achieve this, we couple the parameters of Gaussians with superquadrics.
Given a posed superquadric, centers of Gaussians are determined by uniformly sampling on triangular faces, and poses of Gaussians are determined by face vertices. Inspired by GaMeS [39], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on vertices’ positions. For a given triangular face $V$ and its vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$. The vector $r_{1}$ is aligned with the normal vector of $V$ to ensure Gaussians align with the triangular surface. The vector $r_{2}$ is defined as the vector from the center $m=\text{mean}(v_{1},v_{2},v_{3})$ of the triangle to the first vertex $v_{1}$:
$r_{2}=\frac{v_{1}-m}{||v_{1}-m||}.$ (2)
The last rotation vector $r_{3}$ is obtained by orthonormalizing the vector from the center to the second vertex concerning the existing $r_{1}$ and $r_{2}$ :
@@
where $s_{v}^{2}=c||m-v_{1}||$, $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of Gaussians on each triangular face. These designs eliminate the need to learn the geometry of Gaussians, focusing solely on their texture, thereby enhancing the efficiency of the representation.
Opacity Parameters. We define the total number of hybrid blocks as $M$. However, in practice, a typical scene does not contain exactly $M$ blocks. Therefore, only the meaningful blocks are retained. To achieve this, we introduce a learnable parameter $\tau_{i}$ to represent each block’s opacity. During the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained. Note that Gaussians on the same block share the same $\tau$ for opacity in rasterization.
@@
3.2 Optimizing the Hybrid Representation
This section describes the optimization of the hybrid representation. We found that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks (primitives) accurately. Therefore, we introduce several regularizations and perform two optimization stages: block-level and point-level optimization.
3.2.1 Block-level Optimization

### v4

# Self-supervised Learning of Hybrid Part-aware 3D Representations of 2D Gaussians and Superquadrics

@@ Abstract
Low-level 3D representations, such as point clouds, meshes, NeRFs and 3D Gaussians, are commonly used for modeling 3D objects and scenes. However, cognitive studies indicate that human perception operates at higher levels and interprets 3D environments by decomposing them into meaningful structural parts, rather than low-level elements like points or voxels. Structured geometric decomposition enhances scene interpretability and facilitates downstream tasks requiring component-level manipulation. In this work, we introduce PartGS, a self-supervised part-aware reconstruction framework that integrates 2D Gaussians and superquadrics to parse objects and scenes into an interpretable decomposition, leveraging multi-view image inputs to uncover 3D structural information. Our method jointly optimizes superquadric meshes and Gaussians by coupling their parameters within a hybrid representation. On one hand, superquadrics enable the representation of a wide range of shape primitives, facilitating flexible and meaningful decompositions. On the other hand, 2D Gaussians capture detailed texture and geometric details, ensuring high-fidelity appearance and geometry reconstruction. Operating in a self-supervised manner, our approach demonstrates superior performance compared to state-of-the-art methods across extensive experiments on the DTU, ShapeNet, and real-world datasets.
footnotetext: Corresponding authors
@@ 1 Introduction
Figure 1: Prior works [32, 36] model the scene using primitives, while the proposed method can further model precise geometry details and textures. Here, EMS [32] takes point cloud inputs and reconstructs non-textured primitives.
@@
3D reconstruction from multi-view images is a long-standing challenge in 3D vision and graphics [55, 17, 12]. Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes. However, humans tend to understand 3D scenes as reasonable parts [35]. For instance, when observing a scene, we naturally construct high-level structural information, such as scene graphs, instead of focusing on low-level details like point clouds or voxels. Motivated by it, we propose a part-aware reconstruction framework that decomposes objects or scenes into meaningful shapes or parts, facilitating tasks such as physical simulation, editing, content generation and understanding [18, 28, 58, 53, 54].
Several prior works [72, 37, 29, 32, 33, 56] have explored part-aware reconstruction or 3D decomposition. However, most of them rely heavily on 3D supervision and often struggle to retain fine-grained geometric details, limiting their practicality in real-world scenarios. Recent advances [48, 68] have focused on extending the neural radiance field framework for part-aware reconstruction, building upon its remarkable success in 3D reconstruction from multi-view images. For example, PartNeRF [48] models objects using multiple neural radiance fields. Yet, the intricate composition of implicit fields complicates the learning process, leading to suboptimal rendering quality and inefficient decomposition. Recently, DBW [36] proposes a novel approach that decomposes scenes into block-based representations using superquadric primitives [2], optimizing both their parameters and UV texture maps through rendering loss minimization. While this method demonstrates effective scene decomposition into coherent components, it faces challenges in achieving accurate part-aware reconstruction of both geometry and appearance, as illustrated in Fig. 1.
@@
To address these limitations, we propose PartGS, a self-supervised part-aware hybrid representation that integrates 2D Gaussians [21] and superquadrics [2], to achieve both high-quality texture reconstruction and geometrically accurate part decomposition. There are previous approaches [20, 14, 50] combining mesh reconstruction and Gaussian splatting, where they first reconstruct the mesh and then attach Gaussians to it. In contrast to their primary focus on acquiring the representations, the proposed approach is to achieve part-aware decomposition reconstruction. Our method establishes a coupled optimization framework that simultaneously learns superquadric meshes and Gaussians through parameter sharing, with each part constrained to a single superquadric. The differentiable rendering of Gaussians drives this hybrid representation, leveraging the inherent convexity of superquadrics for qualitative 3D decomposition. The self-supervised training of part-aware reconstruction is built upon the assumption that each part of most objects should be a basic shape and can be represented by a superquadric. Meanwhile, Gaussian splatting, renowned for its superior rendering quality and efficient training, is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality.
In the hybrid representation, 2D Gaussians are distributed on superquadric surfaces to form structured blocks. The pose and shape of Gaussians within each block are determined by their corresponding superquadrics rather than being optimized independently. The parameter space encompasses global controls for superquadric properties (shape, pose, and opacity) and local spherical harmonic coefficients for individual Gaussians. Compared to standard Gaussian splatting [25, 21], which populates the occupancy volume with independently parameterized Gaussians, the coupled representation is more compact and efficient.
@@
During training, the parameters are optimized through rendering loss without additional supervision. However, we notice that image rendering loss often leads to local minima in superquadric shape optimization. To tackle this, we introduce several regularizers to maintain global consistency between the 3D representation and the input 2D information. This strategic implementation allows us to segment parts fully self-supervised, achieving block-level reconstruction. Finally, to better capture irregular shapes, we implement a point-level refinement step that frees 2D Gaussians to deviate from the surface, thereby enhancing geometric fidelity. Extensive experiments show that PartGS, at block-level and point-level, achieves 33.3% and 75.9% improvements in reconstruction accuracy, 3.18 and 16.13 increases in PSNR, and 4X and 3X speedups compared to the state-of-the-art baseline. Our contributions are summarized as follows:
  • We introduce a novel hybrid representation for part-aware 3D reconstruction, combining the strengths of superquadrics and Gaussian splatting to achieve reasonable decomposition and high-quality rendering.
  • We propose several novel regularizers to enforce consistency between 3D decomposition and 2D observations, enabling self-supervised part decomposition.
  • Compared to prior works, the method takes one step forward, by reconstructing both the block-level and point-level part-aware reconstructions, preserving both part segmentation and reconstruction precision.
@@
Figure 2: Overview of our pipeline. PartGS takes multi-view images to learn a parametric hybrid representation of superquadrics and 2D Gaussians. It initializes from random superquadrics and is gradually optimized during training to obtain a block-level reconstruction. Then, we free the constraints of Gaussians to model detailed geometry, achieving point-level reconstruction.
2 Related Work
2.1 Shape Decomposition and Abstraction
Structured shape representation learning decomposes objects or scenes into coherent geometric primitives, facilitating shape understanding and generation [59, 19, 15, 9, 47]. Early approaches like Blocks World [43] and Generalized Cylinders [3] emphasized compact representations. Modern methods typically process 3D inputs (point clouds, meshes, voxels) by decomposing them into primitive ensembles, including cuboids [49, 42], superquadrics [38, 39, 60, 32], and convex shapes [8, 10]. For instance, MonteBoxFinder [42] integrates clustering [11, 51, 62], cuboid representations, and Monte Carlo Tree Search, while EMS [32] adopts a probabilistic approach for superquadric recovery. However, these methods are limited to coarse shape representations. Recent advances in shape abstraction [13, 40, 22, 46, 31] enable more detailed shape representations through flexible primitive deformation.
@@ 2.1 Shape Decomposition and Abstraction
Some studies [48, 36, 68, 1, 16, 70] attempt to create structure-aware 3D representations directly from images. PartNeRF [48] introduces ellipsoid representations within NeRFs, its reliance on multiple implicit fields results in inefficient 3D decomposition. ISCO [1] and DBW [36] use 3D superquadrics for shape decomposition, which enables more meaningful structural separation. However, their simple shape parameters struggle to capture complex geometries, leading to poor geometry and appearance reconstruction. DPA-Net [68] has advanced 3D shape abstraction from sparse views but generates redundant parts and struggles with realistic texture rendering. A concurrent work, GaussianBlock [24], employs SAM [27] to guide superquadric splitting and fusion for 3D decomposition, yet its computational efficiency remains limited, typically requiring several hours per scene. In contrast, our approach accomplishes self-supervised [61, 52] part-aware scene reconstruction through an efficient hybrid representation that simultaneously maintains geometric fidelity and photorealistic rendering quality.
@@ 2.2 Mesh-based Gaussian Splatting
Gaussian splitting (GS) [25] has been rapidly adopted in multiple fields due to its remarkable rendering capability. Several studies [20, 7, 14, 50] attempt to align Gaussians with mesh surfaces for easier editing and animation. Among them, SuGaR [20] uses flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and actual Gaussians. GaMeS [50] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to triangular facets of the mesh. Similarly, Gao et al. [14] proposes a mesh-based GS to achieve large-scale deformation effects on objects. Recently, 2DGS [21] proposes 2D Gaussians for surface modeling and significantly enhances the geometric quality. Inspired by them, we propose a representation that combines 2D Gaussians with superquadrics. A key distinction between our method and previous mesh-based GS approaches is that these methods require first reconstructing the scene’s mesh and then binding Gaussians to the mesh surface, which results in a non-continuous mesh. In contrast, our method directly optimizes the mesh through a rendering loss, enabling part-aware mesh reconstruction.
@@ 3 Method
Given a set of calibrated multi-view images and foreground masks, we aim to learn part-aware 3D representations, a meaningful decomposition of both geometry and appearance. Our approach adopts a two-stage optimization strategy: first, decomposing the object into basic shapes using a mixture of Gaussians and superquadrics at the block-level, followed by refining the decomposition at the point-level to achieve precise geometry. In Sec 3.1, we parameterize the hybrid representation. Sec. 3.2 then elaborates on leveraging this representation, enhanced by novel regularizers and an adaptive control strategy, to achieve self-supervised block-level decomposition. Finally, Sec. 3.3 presents the process for obtaining detailed part-aware results.
3.1 Parametrizing the Hybrid Representation
As shown in Fig. 2, to leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes. This representation retains the superquadric’s ability to parse a 3D scene into distinct parts. Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of prior work [32, 42, 36, 67]. Sharing pose parameters between superquadrics and 2D Gaussians further improves the representation’s efficiency.
The parametric representation is controlled by both primitive and Gaussian parameters, which are optimized simultaneously. Given a 3D scene $\mathcal{S}$, the proposed method decomposes it into multiple hybrid blocks, each consisting of a superquadric with associated Gaussians. Each scene is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$, where $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks. Blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture. These parameters are optimized via differentiable rendering to parse the target scene.
Shape and Scale Parameters. For each hybrid block $\mathcal{B}_{i}$, its geometry is controlled by superquadric parameters [2]. Specifically, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define its shape, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes. Analogous to icosphere vertex positioning, superquadric vertex coordinates are computed by:
@@ 3.1 Parametrizing the Hybrid Representation
where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate. The shape and scale parameters govern block deformation to learn part-aware geometry.
Pose Parameters. The pose of the $i$-th hybrid block is defined by its rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. The vertex position from Eq. 1 is transformed from the local coordinate to world space as: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $j$ indexes the vertices. Previous approaches, including SuGaR [20] and GaMeS [50], position Gaussians on reconstructed meshes with independent pose and shape parameters to enhance appearance modeling. In contrast, our method employs Gaussians to construct a differentiable rendering that bridges images and superquadrics. To achieve this, we couple the parameters of Gaussians with superquadrics.
For a posed superquadric, Gaussian centers are uniformly sampled on triangular faces, with their poses determined by face vertices. Following GaMeS [50], each Gaussian’s rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$ and scaling $\mathrm{S}_{v}$ are computed from vertex positions. Given a triangular face $V$ with vertices $v_{1},v_{2},v_{3}\in\mathbb{R}^{3}$, orthonormal vectors are constructed such that $r_{1}$ aligns with the face normal, $r_{2}$ points from the centroid $m=\text{mean}(v_{1},v_{2},v_{3})$ to $v_{1}$. $r_{3}$ is obtained by orthonormalizing the vector from the center to the second vertex concerning the existing $r_{1}$ and $r_{2}$ :
@@
where $s_{v}^{2}=c||m-v_{1}||$, $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of Gaussians on each triangular face. These designs eliminate the need to learn the geometry of Gaussians, thereby enhancing the efficiency of the representation.
Opacity Parameters. We define the total number of hybrid blocks as $M$. However, a typical scene does not contain exactly $M$ blocks. Therefore, only the meaningful blocks are retained. To achieve this, we introduce a learnable parameter $\tau_{i}$ to represent each block’s opacity. During the optimization, only blocks with $\tau_{i}$ greater than a certain threshold are retained. Note that Gaussians within the same block share the same $\tau$ for opacity in rasterization.
@@
3.2 Block-level Decomposition
This section describes the optimization of the hybrid representation. We observed that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks. Therefore, several regularization terms are introduced to optimize the composition for maximal image formation consistency.
3.2.1 Optimization